OpenStudy (anonymous):

What is the final temperature in the bomb calorimeter if 1.785 g of benzoic acid (HC7H5O2) is combusted in a calorimeter with a heat capacity of 5.02 Kj/degrees C and initial temperature of 24.62 degrees B? The heat of combustion of benzoic acid is -26.42 Kj/g I know that q=m x c x delta T

OpenStudy (aaronq):

\(q=m*C*(T_f-T_i) \rightarrow T_f=\dfrac{q}{m*C}+T_i \)

OpenStudy (chmvijay):

q=-26.42 Kj/g C=5.02 Kj/ m=1.785 g Ti=24.62 °C find Tf